OpenStudy (anonymous):

if f(x) = x - 7, find f^-1 (3). Could someone teach me how to solve this? My teacher didn't go over this and I can't find this on my text book.

OpenStudy (amistre64):

inverse functions, in general, swap xs and ys

OpenStudy (amistre64):

y = x-7 the inverse is found by swapping xs and ys x = y-7 then "resolve" the equation for y again

OpenStudy (amistre64):

or you can just go ahead with the inverse setup and place in the 3 for x and solve 3 = y-7

OpenStudy (anonymous):

3 = y - 7? aren't you supposed to substitute y for 3?

OpenStudy (amistre64):

no; the inverse function is still a function of x, so the proper thing to do would be to solve for y again by getting it by its little lonesome

OpenStudy (amistre64):

at any rate, all xs are 3s in this case

OpenStudy (anonymous):

wait, lemme show you mine lol

OpenStudy (anonymous):

AMG i forgot how i did this.

OpenStudy (amistre64):

jsut start at the beginning :)

OpenStudy (anonymous):

f(x) = x - 7 f(3) = 3 - 7 f(3) = -4 f^-1 (x) = 4 so uhh.. lol

OpenStudy (amistre64):

an inverse is not looking for the value of the original "x"; its looking for the value that creates a 3 in the original. f(x) = y = x-7 ; when x=10, y=3 swap out variables: x = y - 7 resolve: f^-1(x) = y = x + 7 when x=3, y=10

OpenStudy (amistre64):

an inverse undoes the original to get you back to x again

OpenStudy (anonymous):

when x = 10, y = 3 Swap out the variables: f(x) = x - 7 y = x - 7 3 = 10 - 7 3 = 3 Resolve: f^-1(x) = x + 7 y = x + 7 I'm sorry I still can't understand. I'll try to do some research.

OpenStudy (amistre64):

alot of the confusion might be in using "x" in both instances.

OpenStudy (amistre64):

\[y = f(x)\to\ f^{-1}(y)=x\]a function gets us from x to y; the inverse gets us from y back to x again

OpenStudy (amistre64):

the issue is, that by convention we use x as an independant variable. By defining an inverse as a function of x, we create a more confusing learning environment

OpenStudy (anonymous):

a my god i can't find anything over the internet, what the heck did this teacher give us >.<

OpenStudy (anonymous):

@amistre64 hey! i think i figured it out, how about this? f(x) = x - 7 f(10) = 10 - 7 f(10) = 3 f^-1(3) = 10

OpenStudy (amistre64):

that is actually a good read.